Let 𝑓(π‘₯) = ⌊π‘₯βŒ‹ + βŒŠβˆ’π‘₯βŒ‹, where ⌊π‘₯βŒ‹ is the greatest integer less than or equal to π‘₯.


(π‘Ž) For what values of π‘Ž, does limπ‘₯β†’π‘Ž


𝑓(π‘₯)exist?


(𝑏) At what numbers is 𝑓 discontinuous?

1
Expert's answer
2022-03-25T15:29:33-0400

(a) Let "a\u2208\u211d" such that "a\u2209\u2124" .

let "\lfloor{a}\rfloor=m"

"=> \lfloor{-a}\rfloor=-(m+1)"


There exist "\u03b5>0" such that "\u2200 r\u2208(a-\u03b5,a+\u03b5)"

"=> \lfloor{r}\rfloor=m"

and "\lfloor{-r}\rfloor=-(m+1)"


Thus,

"\lim\limits_{x\rarr a^{+}}f(x)=\lim\limits_{x\rarr a^{-}} f(x)=\lim\limits_{x\rarr a}f(x)"

"=\lfloor{r}\rfloor+\lfloor{-r}\rfloor=m-(m+1)=-1"



Suppose "a\u2208\u2124" . There exist "\u03b5>0" such that "\u2200 k\u2208(a-\u03b5, a)"

"=> \lfloor{k}\rfloor=a-1"

and "\lfloor{-k}\rfloor=-a"


"\lim\limits_{x\rarr a^{-}}f(x)=\lfloor{k}\rfloor+\lfloor{-k}\rfloor=(a-1)-a=-1"


Also, there exists "\u03b5>0" such that "\u2200 k\u2208(a, a+\u03b5)"

"=> \lfloor{k}\rfloor=a"

and "\lfloor{-k}\rfloor=-(a+1)"


"\lim\limits_{x\rarr a^{+}}f(x)=\lfloor{k}\rfloor+\lfloor{-k}\rfloor=a-(a+1)=-1"


"\lim\limits_{x\rarr a^{+}}f(x)=\lim\limits_{x\rarr a^{-}} f(x)=\lim\limits_{x\rarr a}f(x)=-1"


Hence, "\lim\limits_{x\rarr a}f(x)" exists for all "a\u2208\u211d"






(b) Let "x\u2208\u211d" such that "x\u2209\u2124"

let "\lfloor{x}\rfloor=m"

"=>\lfloor{-x}\rfloor=-(m+1)"


"f(x)=\lfloor{x}\rfloor+\lfloor{-x}\rfloor=m-(m+1)=-1"


We established in (a) above that "\lim\limits_{x\rarr a}f(x)=-1" , where a is any non-integer


Thus, "\lim\limits_{x\rarr a}f(x)=f(x)" .


In addition, "f(x)" is defined "\u2200 x\u2208\u211d"


Thus, f is continuous at any non-integer point



Suppose "x\u2208\u2124" .

"f(x)=\lfloor{x}\rfloor+\lfloor{-x}\rfloor=x-x=0"


We established in (a) above that "\lim\limits_{x\rarr a}f(x)=-1" , where a is any integer


Thus, "\lim\limits_{x\rarr a}f(x)≠f(x)"

Hence, f is discontinuous at any integer point.


f is discontinuous "\u2200x\u2208\u2124"